OpenStudy (anonymous):

What are the solutions? 16x^2+9=0

OpenStudy (anonymous):

I see two perfect squares here... 16 and 9. what are their components?

OpenStudy (anonymous):

3 and 4?

OpenStudy (anonymous):

oh wait that says +9...

OpenStudy (anonymous):

when in doubt use the quadratic formula i guess... \[\frac{ \pm \sqrt{-4(16)(9)} }{ 2(16) }\]

OpenStudy (anonymous):

i see something wrong here...

OpenStudy (anonymous):

\[-\frac{ 3 }{ 4 }i,\frac{ 3 }{ 4 }i\] \[-\frac{ 4 }{ 3 }i,\frac{ 4 }{ 3 }i\] \[-\frac{ 3 }{ 4 },\frac{ 3 }{ 4 }\] \[-\frac{ 9 }{ 16 }i,\frac{ 9 }{ 16 }i\]

OpenStudy (anonymous):

ohh its imaginary... thats why

OpenStudy (anonymous):

what is the (i) for in these? i dont understand it and it gets me confused sometimes

OpenStudy (anonymous):

ok so you see that quadratic formula that I did up there? I got the square root of a negative which isn't real. Those i's in your answer however indicate that the answer is imaginary. Solving the quadratic equation by taking the negative out and putting an i. So, you get: \[\pm \frac{ 3i }{ 4 }\]

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

can u help me with some more questions? @piglet9

OpenStudy (anonymous):

I can try just mention me and ill try to get to em